OpenStudy (anonymous):

Write the complex number in rectangular form 7(cos 120° + i sin 120°)

OpenStudy (john_es):

I assume rectangular notation is a+b*i, then you should only calculate the value of trigonometric functions, \[\cos 120=-1/2\\ \sin120=\sqrt{3}/2\] and then multiply by 7.